How many grams of Ca metal are needed to react with 15.3 liters of N2 gas at STP?
3 Ca(s) + N2(g) → Ca3N2(s)
1
Expert's answer
2020-10-26T13:29:26-0400

n (N2) ="V\/V"m n (N2)=15,3/22,4=0,683 mol n(Ca)=3*0,683=2 mol

m( Ca)=n*M=2*40=80 g.